Reports 1-1 of 1 Clear search Modify search
IOO (General)
kenta.tanaka - 7:36 Wednesday 23 November 2022 (23008) Print this report
IMC cannot lock

IMC suddenly got down at 4:45 JST and has not returned since then to the present.

### What I did

  • We were using Yarm to measure the sensing matrix. When we came to Mozumo last night around 7 pm, Yarm was already locked, and it stayed locked until 4 am when it got down.
  • When preparing the measurement files for ITMY, we noticed that Yarm and IMC were down. At that time, we thought they would be back up soon. However, IMC did not come back.
  • I checked the alignment at the time of the IMC down with their oplevs and found no change before or after the fall. Also, as far as I could see with the transmitted light camera, it appeared to be flashed by TEM00. In other words, it does not seem that the lock does not return due to bad alignment. (Fig. 1)
  • According to the IMC transmitted camera image, IMC locked at the moment but did not seem to be able to maintain the lock, perhaps because there was not enough gain after that.
  • When I tried to lock manually, I found that after turning on the SLOW path boost filter, raising IN1GAIN from the nominal 4dB to about 12dB improved the lock duration to a few tens seconds, but then adding even one boost filter in the common path caused IMC down.
  • In addition, when we looked at the magnitude of the error signal, we noticed that it was clearly smaller than before (Before: Fig. 2, Now: Fig.3). On the other hand, the amount of light in the RFPD had not changed.(Fig. 4)
  • I think we need to go down into the mine to find out why IMC is not locking up any more.
Images attached to this report
Comments to this report:
shinji.miyoki - 22:17 Wednesday 23 November 2022 (23010) Print this report

[miyoki. uehara, tanaka, hirose, yokozawa]


Two Keysight N5181B signal generators(SGs) were broken. One is 13.78MHz for IMC lock, The other is 5.6MHz for IFO.

One for 13.78Mhz was replaced with an unused same-model SG that was put in the front clean room of the PSL room. One for 5.6MHz was replaced with the E8663D.

Because the setting for SG 13.78MHz was recovered, the IMC lock came back. While the output of SG 5.6MHz was set at -60 dbm, not 13.3 dbm to avoid some trouble by us. So we need 5.6MHz signal recovery tomorrow. 



As Tanaka-kun pointed out, the error signal of IMC was small. The pp-amplitude is only 0.008 that should be 0.8. And, the control UGF is only 1kHz! Also, no failure of cable connection and optical condition were reported. So, the most suspicious trouble was "no modulation or no demodulation" process.

Firstly, we checked the RF signal at the modulation EOM input point, and the RF signal exists by an oscilloscope, so no problem.

Secondly, we checked the error signal at the just output of the mixer in the LSC?IMC? rack, then a very weak PDH-signal was confirmed. 

3rdly, we checked the RF signal from the IMC-REFL-PD at the mixer point by using a spectrum analyzer, then we found no signal at 13.78MHz but 17.4MHz! So we suspected the RF signal itself has trouble.

4thly, we checked the 13.78Mhz RF signal itself from the distributor/N5181B in the rack, then we found it generated not 13.78MHz but 17.4MHz. Also we found "Error 508: Synthesizer unlock". We also found "Error 512: Reference unlock" in SG for 5.6MHz.

According to the web manual, it seems to require an" A3 RF assembly change". Because we don't use references for these SGs, we should go "Post-Repair Procedures and OperationalPerformanceVerification" step. The latest firmware installation is recommended.

However, the broken conditions are different with each other. The SG for 13.7Mhz actually generated 17.4MHz signal according to a spectrum analyzer. 17.8MHz itself seems to be stable. However, The SG for 5.6MHz generated the 5.6MHz "before rebooting". After rebooting, this SG generated ~ 2MHz "peak-fluctuating signal", not 5.6MHz. These same strange performance tendency seems to be originated these errors.

We tried self-test for both SGs, and both showed 5~6 errors and the message as the attached 4th photo.


Temporal measures

The SG for 13.78Mhz was replaced with the unused same-model SG that was put in the front clean room of the PSL room (1st photo). This unused one passed its self-test and showed no errors. So we removed the two broken SGs from the rack and reinstall this SG, and resumed the RF frequency and set the amplitude at -60dbm. After reconnecting cables to this SG, we turned on the RF out, and increased the amplitude to 7.4dbm gradually. After this, IMC can be locked. The alignment was not performed, just a lock.

While the SG for 5.6mHz was put on the "sunoko" on the floor beside the rack, and we resumed the frequency (5.6246235MHz) at the amplitude of -60dbm(2nd photo before putting on sunoko). We connected this SG and Wenzel distributor with ~ 3m RF cable, then set the RF out on. It should be 13.3dbm according to the original setting. To avoid any too much voltage applying to the distributor, we left it at -60dbm.



  1. IMC alignment. UGF check.
  2. Recovery of 5.6MHz signals and IFO.
  3. Urgent preparation for SGs if we have no more SGs, or better SG usage optimization according to each requirement.
  4. Serial number record changes for these SGs replacement?
Images attached to this comment
shinji.miyoki - 23:49 Thursday 24 November 2022 (23025) Print this report

Not 13.3 dBm but 10.0 dBm is selected for 5.6MHz signal.


shinji.miyoki - 7:49 Friday 25 November 2022 (23028) Print this report

According to this klog#8247, a proper input amplitude value for the Wenzel harmonics generator is desirable to obtain enough amplitude for some harmonics.

Then klog#8422 was tried (detail measurement condition) and 13.3dBm seemed to be proper to obtain "specified values by Wenzel Inc." for some harmonics.

The data that was provided by Wenzel Inc. is here.

shinji.miyoki - 9:50 Friday 25 November 2022 (23032) Print this report

Amplitude measurement for the Wenzel harmonics generator was done as reported, while the phase measurement was not done. High multiple harmonics could have phase delay changes according to their amplitude.

shinji.miyoki - 20:48 Friday 25 November 2022 (23038) Print this report

I swapped the SG. So now N5181B for 5.6 MHz:13.3dBm, and E8663D for 13.78MHz:7.4dBm.

Both SGs are now controllable on the web. IP information is restricted only to Ushiba-kun and Yamamoto-kun, and Oshino-kun.

shinji.miyoki - 21:01 Monday 28 November 2022 (23061) Print this report

I updated the firmware for these broken SGs.

When I connected the external trigger signal (10MHz) from one more oscillator, the errors disappeared(Before updating, errors were shown even if the external signal was adopted). However, the setting freq. and the actual output freq. still had discrepancies.

While the internal reference still showed errors:512. In addition, the 10MHz signal output on the rear panel of one of SGs was also broken. 

So we need repair for them.

Search Help